# Traffic Scenario Logic: A Spatial-Temporal Logic for Modeling and Reasoning of Urban Traffic Scenarios

### v1

@@ 4.2 Rules of Highway TSL
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\\ \to\mdwhtcircle(\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\lor\mathsf{lonr}(c_{1},c_{2},\mathsf{cover}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\\ \to\mdwhtcircle(\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\lor\mathsf{lonr}(c_{1},c_{2},\mathsf{cover}))).$
  5. 5.
    The occupied lanes must be continuous, meaning that the lanes occupied by a vehicle must be one or more adjacent lanes. The logical representation of this rule is slightly complex. First, we introduce an auxiliary predicate $\mathsf{cleft/2}$, representing one lane being to the left of another lane but not necessarily adjacent. This can be achieved by defining it recursively as follows:
    $\mdwhtsquare(\forall l_{1}\forall l_{2}\,\mathsf{left}(l_{1},l_{2})\to\mathsf{cleft}(l_{1},l_{2})),$
    $\mdwhtsquare(\forall l_{1}\forall l_{2}\forall l_{3}\,\mathsf{left}(l_{2},l_{3})\land\mathsf{cleft}(l_{1},l_{2})\to\mathsf{cleft}(l_{1},l_{3})).$
    Therefore, the rule stating that the occupied lanes must be continuous can be expressed as:
    $\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\forall l_{3}\,\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\\ \land\mathsf{cleft}(l_{1},l_{3})\land\mathsf{cleft}(l_{3},l_{2})\to\mathsf{on}(c,l_{3})),$
    which means any lane between two lanes occupied by a vehicle must also be occupied by that vehicle.
  6. 6.
@@
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\\ \land\lnot(l_{1}=l_{2})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{1})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{2}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\lnot\mathsf{on}(c,l_{1})\land\lnot\mathsf{on}(c,l_{2})\\ \land\lnot(l_{1}=l_{2})\land\mdwhtcircle\mathsf{on}(c,l_{1})\land\mdwhtcircle\mathsf{on}(c,l_{2}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\mathsf{on}(c,l_{1})\land\lnot\mathsf{on}(c,l_{2})\\ \land\lnot(l_{1}=l_{2})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{1})\land\mdwhtcircle\mathsf{on}(c,l_{2}))).$
@@
    $\mdwhtsquare\forall c\forall l_{1}\forall l_{2}\forall l_{3}\,\lnot(\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\land\mathsf{on}(c,l_{3})\\ \land\lnot(l_{1}=l_{2})\land\lnot(l_{1}=l_{3})\land\lnot(l_{2}=l_{3})).$
  2. 2.
@@
    $\mdwhtsquare\forall c_{1}\forall c_{2}\forall l\,\lnot(\mathsf{on}(c_{1},l)\land\mathsf{on}(c_{2},l)\\ \land\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\land\lnot(c_{1}=c_{2})).$
A realistic scenario must satisfy all the aforementioned rules. Consequently, by utilizing these rules, we can validate the plausibility of a scenario, rendering the reasoning of scenarios under TSL executable.
5 Traffic Scenario Logic for Urban
@@ 5.1 Formal Representation of Urban Road Network
A structured urban road network can be described by an OpenDRIVE map, as shown in Figure 2. Considering two lanes on the ground, their relationships can be categorized as follows:
@@
    They have an intersection point, but are not connected at this point;
@@
    They are connected at a connection point, where vehicles can switch to another lane;
@@
    They have an overlap segment in which vehicles travel in the same direction. In this case, it can be considered as one lane;
@@
    They have an overlap segment in which vehicles travel in opposite directions;
@@
All we need is to extend the TSL for intersections, connections, and overlapping in opposite directions. The rest of the cases can reuse the logical representation defined in Section 4.
@@
    Denote the intersection point as $p^{\mathsf{x}}$. Then, for any vehicle $c$ on these two intersecting lanes, there are essentially three positional relationships with respect to this intersection point: yet to pass, currently passing, and already passed. We use the predicate $\mathsf{lonpr/3}$ to represent these relationships, corresponding to the predicate $\mathsf{lonr/3}$ for vehicles’ relationships. The same predicate can also be used for the connecting points and endpoints of overlapping segments below.
@@
    Two or more lanes can connect at their endpoints. We denote the connecting points as $p^{\mathsf{c}}$. Depending on the direction of travel of vehicles on the lanes, there exists a successor relationship $\succ_{\mathsf{c}}$ between the connecting points and the lanes.
@@
    Overlapping in opposite directions
    When examining the overlapping segment between points $p^{\mathsf{os}}$ and $p^{\mathsf{oe}}$, as shown in Figure 3, we can designate one of the two overlapping lanes’ direction as the reference direction, i.e. from $p^{\mathsf{os}}$ to $p^{\mathsf{oe}}$. Another longitudinal positional relationship, denoted as $\mathsf{lonro/3}$, exists between the vehicles in the overlapping segment, and:
@@
$p^{\mathsf{os}}$$p^{\mathsf{oe}}$
Figure 3: A overlapping segment of two lanes with opposite directions
Besides, when there are multiple points (including intersection points, connection points, and endpoints of overlapping segments) on a lane $l$, there also exists a successor relationship $\succ_{\mathsf{p}}^{l}$ between points according to the direction of vehicle travel on the lane.
@@
    $\mathbb{O}\subseteq\mathbb{P}^{\mathsf{OS}}\times\mathbb{P}^{\mathsf{OE}}$ represents the pairs $(p^{\mathsf{os}},p^{\mathsf{oe}})$ for each overlapping segment.
@@
2(b) shows the topology abstraction of a typical T-intersection based on Figure 3. There are 12 lanes in it, with 6 connection points (marked by blue diamonds) and 3 intersection points (marked by red circled times).
@@ 5.2 Abstract Model
    $\mathscr{P}:\mathbb{C}\times\mathbb{P}\mapsto\mathbb{D}$ is a mapping from pairs of vehicles and points to the set of longitudinal positional relationships, as explained in subsection 5.1;
    $\mathscr{Q}:\mathbb{C}\times\mathbb{C}\mapsto\mathbb{D}$ is a mapping to define the temporary longitudinal positional relationship $\mathsf{lonro/3}$ on the overlapping segments.
@@
[Scenario] A (urban) scenario of is a sextuple $\mathcal{S}=(\mathbb{C},\mathbb{L},\mathbb{D},\mathbb{P},T,\langle\mathfrak{s}_{0},\mathfrak{s}_{1},\dots,\mathfrak{s}_{T-1}\rangle)$, where:
@@
The definition of the tail of a scenario is similar to subsection 4.1 and will not be repeated.
@@
[The Semantics of Traffic Scenario Logic for Urban] Given the road network $\mathcal{M}$ and the logical scenario $\mathcal{S}$, the satisfaction of formulas defined in subsection 5.2 follows the following semantics:
@@ 5.3 Rules of Urban TSL
For urban traffic scenarios, the rules described in subsection 4.2 regarding vehicles within the same road still apply and will not be reiterated here. Additionally, the following rules need to be added to handle situations across lanes:
  1. 1.
    A vehicle can only be located on one road. Firstly, we introduce a auxiliary predicate $\mathsf{cbelong/2}$ to represent the belonging relationship between a vehicle and a road:
    $\mdwhtsquare(\forall c\forall l\forall\mathcal{R}\,\mathsf{belong}(l,\mathcal{R})\land\mathsf{on}(c,l)\to\mathsf{cbelong}(c,\mathcal{R})).$
    Then we have
    $\mdwhtsquare(\forall c\,\forall\mathcal{R}_{1}\forall\mathcal{R}_{2}\,\lnot(\mathsf{cbelong}(c,\mathcal{R}_{1})\land\\ \mathsf{cbelong}(c,\mathcal{R}_{2})\land\lnot(\mathcal{R}_{1}=\mathcal{R}_{2}))).$
  2. 2.
@@
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{behind})\to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{behind})\\ \lor\mathsf{lonpr}(c,p,\mathsf{cover})\lor\mathsf{lonpr}(c,p,\mathsf{none}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{cover})\to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{cover})\\ \lor\mathsf{lonpr}(c,p,\mathsf{ahead})\lor\mathsf{lonpr}(c,p,\mathsf{none}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{ahead})\\ \to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{ahead})\lor\mathsf{lonpr}(c,p,\mathsf{none}))).$
  3. 3.
@@
    $\mdwhtsquare(\forall c\forall p\forall d_{1}\forall d_{2}\,\lnot(\mathsf{lonpr}(c,p,d_{1})\\ \land\mathsf{lonpr}(c,p,d_{2})\land\lnot(d_{1}=d_{2}))),$
  4. 4.
@@
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p\,\lnot(\mathsf{lonpr}(c_{1},p,\mathsf{cover})\\ \land\mathsf{lonpr}(c_{2},p,\mathsf{cover})\land\lnot(c_{1}=c_{2}))).$
  5. 5.
@@
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\forall l_{1}\,\mathsf{lonpr}(c,p,\mathsf{cover})\land\mathsf{p_{c}}(p)\land\mathsf{on}(c,l_{1})\to\\ \mdwhtcircle((\mathsf{lonpr}(c,p,\mathsf{cover})\land\mathsf{on}(c,l_{1}))\lor\\ \exists l_{2}(\mathsf{lonpr}(c,p,\mathsf{ahead})\land\mathsf{on}(c,l_{2})\land\mathsf{succl}(p,l_{2})))).$
  6. 6.
    Rules for overlapping segments following the discussion in subsection 5.1. Firstly, we define predicates $\mathsf{fwdover/3}$ and $\mathsf{rvsover/3}$ to indicate the vehicle driving along the reference direction and against the reference direction on an overlapping segment, respectively.
    $\mdwhtsquare(\forall c\forall p_{1}\forall p_{2}\forall l\,\mathsf{lonpr}(c,p_{1},\mathsf{ahead})\land\mathsf{lonpr}(c,p_{2},\mathsf{behind})\\ \land\mathsf{on}(c,l)\land\mathsf{pon}(p_{1},l)\land\mathsf{pon}(p_{2},l)\\ \land\mathsf{overlap}(p_{1},p_{2})\to\mathsf{fwdover}(c,p_{1},p_{2})),$
    $\mdwhtsquare(\forall c\forall p_{1}\forall p_{2}\forall l\,\mathsf{lonpr}(c,p_{1},\mathsf{behind})\land\mathsf{lonpr}(c,p_{2},\mathsf{ahead})\\ \land\mathsf{on}(c,l)\land\mathsf{pon}(p_{1},l)\land\mathsf{pon}(p_{2},l)\\ \land\mathsf{overlap}(p_{1},p_{2})\to\mathsf{rvsover}(c,p_{1},p_{2})).$
    Then we have the formulae about $\mathsf{lonro/3}$ as mentioned above:
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\forall d\,\\ \mathsf{fwdover}(c_{1},p_{1},p_{2})\land\mathsf{fwdover}(c_{2},p_{1},p_{2})\\ \land\mathsf{lonr}(c_{1},c_{2},d)\to\mathsf{lonro}(c_{1},c_{2},d)),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\\ \land\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{behind})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\\ \land\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{cover})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\\ \land\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{ahead})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{fwdover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\\ \to(\exists d\,\mathsf{lonro}(c_{1},c_{2},d)\land\lnot(d=\mathsf{none}))).$
  7. 7.
    All longitudinal positional relationships, including those between vehicles and vehicles, vehicles and points, and those on overlapping segments, follow similar rules as presented in subsection 4.2, which will not be reiterated here.

### v3

@@ 4. PR4
  5. PR5
    The occupied lanes by a vehicle must be one or more adjacent lanes. We utilize an auxiliary predicate $\mathsf{cleft/2}$ (AP1) here for simplicity, representing one lane being to the left of another lane but not necessarily adjacent.
  6. PR6
@@ 1. TR1
  2. TR2
@@
A realistic scenario and its tails (because a tail of a realistic scenario is also a realistic scenario) must satisfy all the aforementioned rules. Consequently, by utilizing these rules, we can validate the plausibility of a scenario, rendering the reasoning of scenarios under TSL executable.
4 Urban Traffic Scenario Logic
@@ 4.1 Formal Representation of Urban Road Network
A structured urban road network consists of lanes. When considering two lanes, their relationships can be categorized as follows:
@@
    They have an intersection point, but are not connected at this point (Figure 1(a));
@@
    They are connected at a connection point, where vehicles can switch to another lane (Figure 1(b));
@@
    They have an overlap segment in which vehicles travel in the same direction. In this case, we can merge the overlap segment into a single lane;
@@
    They have an overlap segment in which vehicles travel in opposite directions (Figure 1(c));
@@
All we need is to extend the TSL for intersections, connections, and overlaps in opposite directions. The rest of the cases can reuse Highway TSL.
@@
    Denote the intersection point as $p^{\mathsf{x}}$. Then, for any vehicle $c$ on these two intersecting lanes, there are essentially three positional relationships with respect to this intersection point: yet to pass, currently passing, and already passed. We use the predicate $\mathsf{lonpr/3}$ to represent these relationships, corresponding to the predicate $\mathsf{lonr/3}$ for vehicles’ relationships. The same predicate can also be used for the connection points and endpoints of overlap segments below.
@@
    Two or more lanes can connect at their endpoints. We denote the connection points as $p^{\mathsf{c}}$. Depending on the direction of travel of vehicles on the lanes, there exists a successor relationship $\succ_{\mathsf{c}}$ between the connection points and the lanes.
@@
    Overlap in opposite directions
    When examining the overlap segment between points $p^{\mathsf{os}}$ and $p^{\mathsf{oe}}$, as shown in Figure 1(c), we can designate one of the two overlapping lanes’ direction as the reference direction, i.e., from $p^{\mathsf{os}}$ to $p^{\mathsf{oe}}$. Another longitudinal positional relationship, denoted as $\mathsf{lonro/3}$, exists between the vehicles in the overlap segment, and:
@@
(a) Intersection
(b) Connection
(c) Overlap
Figure 1: Spectial points on urban road network
Besides, when there are multiple points (including intersection points, connection points, and endpoints of overlap segments) on a lane $l$, there also exists a successor relationship $\succ_{\mathsf{p}}^{l}$ between points according to the direction of vehicle travel on the lane.
@@
    $\mathbb{O}\subseteq\mathbb{P}^{\mathsf{OS}}\times\mathbb{P}^{\mathsf{OE}}$ represents the pairs $(p^{\mathsf{os}},p^{\mathsf{oe}})$ for each overlap segment.
@@
We proposed an example of this abstraction in the Technical Appendix.
Theorem 1
For a structured 2D OpenDRIVE road network, there exists a representation in the form defined by Figure 1.
Proof.
See the discussion above. ∎
@@ 4.2 Abstract Model
    $\mathscr{P}:\mathbb{C}\times\mathbb{P}\mapsto\mathbb{D}$ is a mapping from pairs of vehicles and points to the set of longitudinal positional relationships, as explained in Section 4.1;
    $\mathscr{Q}:\mathbb{C}\times\mathbb{C}\mapsto\mathbb{D}$ is a mapping to define the temporary longitudinal positional relationship $\mathsf{lonro/3}$ on the overlap segments.
@@
[Scenario] A (urban) scenario is a sextuple $\mathcal{S}=(\mathbb{C},\mathbb{L},\mathbb{D},\mathbb{P},T,\langle\mathfrak{s}_{0},\mathfrak{s}_{1},\dots,\mathfrak{s}_{T-1}\rangle)$, where:
@@
The definition of the $i$-tail of a scenario is similar to Section 3.1 and will not be repeated.
Theorem 2
For a pedestrian-free traffic scenario on road network $\mathcal{M}$, there exists a representation in the form defined by Section 4.2.
Proof.
Since the longitudinal relationship as defined by Section 3.1 encompasses all situations, the mappings $\mathscr{L}$, $\mathscr{D}$, $\mathscr{P}$, and $\mathscr{Q}$ are guaranteed to exist and be unique for each scene. Consequently, we establish the existence of a representation $\mathcal{S}$ for each pedestrian-free traffic scenario. ∎
@@
[The Semantics of Urban TSL] Given the road network $\mathcal{M}$ and the logical scenario $\mathcal{S}$, the satisfaction of formulas defined in Theorem 2 follows the following semantics:
@@
Rule TSL Formula
PR8 $\mdwhtsquare(\forall c\,\forall\mathcal{R}_{1}\forall\mathcal{R}_{2}\,\lnot(\mathsf{cbelong}(c,\mathcal{R}_{1})\land\mathsf{cbelong}(c,\mathcal{R}_{2})\land\lnot(\mathcal{R}_{1}=\mathcal{R}_{2})))$.
PR9 $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{behind})\to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{behind})\lor\mathsf{lonpr}(c,p,\mathsf{cover})\lor\mathsf{lonpr}(c,p,\mathsf{none})))$,
$\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{cover})\to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{cover})\lor\mathsf{lonpr}(c,p,\mathsf{ahead})\lor\mathsf{lonpr}(c,p,\mathsf{none})))$,
$\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{ahead})\to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{ahead})\lor\mathsf{lonpr}(c,p,\mathsf{none})))$.
PR10 $\mdwhtsquare(\forall c\forall p\forall d_{1}\forall d_{2}\,\lnot(\mathsf{lonpr}(c,p,d_{1})\land\mathsf{lonpr}(c,p,d_{2})\land\lnot(d_{1}=d_{2})))$,
PR11 $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p\,\lnot(\mathsf{lonpr}(c_{1},p,\mathsf{cover})\land\mathsf{lonpr}(c_{2},p,\mathsf{cover})\land\lnot(c_{1}=c_{2})))$.
PR12 $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\forall l_{1}\,\mathsf{lonpr}(c,p,\mathsf{cover})\land\mathsf{p_{c}}(p)\land\mathsf{on}(c,l_{1})\to$
$\quad\quad\mdwhtcircle((\mathsf{lonpr}(c,p,\mathsf{cover})\land\mathsf{on}(c,l_{1}))\lor\exists l_{2}(\mathsf{lonpr}(c,p,\mathsf{ahead})\land\mathsf{on}(c,l_{2})\land\mathsf{succl}(p,l_{2}))))$.
PR13 $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\forall d\,\mathsf{fwdover}(c_{1},p_{1},p_{2})\land\mathsf{fwdover}(c_{2},p_{1},p_{2})\land\mathsf{lonr}(c_{1},c_{2},d)\to\mathsf{lonro}(c_{1},c_{2},d))$,
$\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\land\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{behind}))$,
$\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\land\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{cover}))$,
$\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\land\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{ahead}))$,
$\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{fwdover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\to(\exists d\,\mathsf{lonro}(c_{1},c_{2},d)\land\lnot(d=\mathsf{none})))$.
AP2 $\forall c\forall l\forall\mathcal{R}\,\mathsf{belong}(l,\mathcal{R})\land\mathsf{on}(c,l)\to\mathsf{cbelong}(c,\mathcal{R})$.
AP3 $\forall c\forall p_{1}\forall p_{2}\forall l\,\mathsf{lonpr}(c,p_{1},\mathsf{ahead})\!\land\!\mathsf{lonpr}(c,p_{2},\mathsf{behind})\!\land\!\mathsf{on}(c,l)\!\land\!\mathsf{pon}(p_{1},l)\!\land\!\mathsf{pon}(p_{2},l)\!\land\!\mathsf{overlap}(p_{1},p_{2})\!\to\!\mathsf{fwdover}(c,p_{1},p_{2})$.
AP4 $\forall c\forall p_{1}\forall p_{2}\forall l\,\mathsf{lonpr}(c,p_{1},\mathsf{behind})\!\land\!\mathsf{lonpr}(c,p_{2},\mathsf{ahead})\!\land\!\mathsf{on}(c,l)\!\land\!\mathsf{pon}(p_{1},l)\!\land\!\mathsf{pon}(p_{2},l)\!\land\!\mathsf{overlap}(p_{1},p_{2})\!\to\!\mathsf{rvsover}(c,p_{1},p_{2})$.
Table 2: The additional rules and auxiliary predicates used in Urban TSL
@@ 4.3 Rules of Urban TSL
For urban traffic scenarios, the rules described in Section 3.2 regarding vehicles within the same road still apply and will not be reiterated here. Additionally, the following rules are added to handle situations across roads, whose formulae are shown in Table 2.
  1. PR8
    A vehicle can only be located on one road. We introduce an auxiliary predicate $\mathsf{cbelong/2}$ (AP2) to represent the belonging relationship between a vehicle and a road.
  2. PR9
@@
  3. PR10
@@
  4. PR11
@@
  5. PR12
@@
  6. PR13
    Rules for overlap segments following the discussion in Section 4.1. Auxiliary predicates $\mathsf{fwdover/3}$ (AP3) and $\mathsf{rvsover/3}$ (AP4) are imported for this rule to indicate the vehicle driving along/against the reference direction on an overlap segment, respectively.
  7. PR14
    All longitudinal positional relationships, including those between vehicles and vehicles, vehicles and points, and those on overlap segments, follow similar rules as presented in Section 3.2, which will not be reiterated here.